OpenStudy (anonymous):

solve f(x)=lx-3l-5 find f(1)-f(5)

OpenStudy (anonymous):

\[f(1)-f(5)=\left| 1-3 \right|-5-(\left| 5-3 \right|-5)=\left| -2 \right|-5-\left| 2 \right|+5=0\]

OpenStudy (anonymous):

thanks!